ABC Inc. recently issued $1,000 par bonds at a 5.25% coupon rate. The bonds have 15 years to maturity and current price of the bond is $850. If the call price is $1,050 and the bond can be called in 10 years, what is the yield to call? Assume semi-annual compounding.

Note: Convert your answer to percentage and round off to two decimal points. Do not enter % in the answer box.
Answer #1

Here

YTC = Interest +(Face value -call price/n) / (Face value + call price/2)

N = no of year in which bond can be called = 10 x 2 = 20

Interest = 1000 x 5.25%/2  = 26.25$

Face vale = 1000$

Call price = 1050 $

Thus Yield to call = 26.25 + (1000-1050)/20 / (1000+1050)/2

=26.25+ ( -50/20) / (2050/2)

=26.25 -2.5 / 1025

=23.75/1025

=0.023171

i.e 2.3171%

Annual YTC = 2.3171% x 2

=4.63%